# Metastability in the Dissipative Quantum Rabi Model

### v1

@@ II.3 Stability under strong perturbations
From the cumulant analysis, we have shown that strong quantum fluctuations exist in the SMP. This indicates that the parity-breaking solutions in the SMP can be only stable for a finite time before a large perturbation drives the system away. We thus refer to these parity-breaking solutions as the metastable solutions and label the phase as the superradiant metastable phase (SMP).
Figure 3: The evolution of the steady-state solution in the NP (a) and SMP (b) under strong perturbations within the mean-field framework. Phenomenological interpretations are shown in the top panels. For the NP, we consider a quench $\langle\hat{x}\rangle=r\cos 2\theta$, $\langle\hat{p}\rangle=r\sin 2\theta$ of the parity-preserving solution. For the SMP, we consider a quench $\langle\hat{\sigma}_{x}\rangle=\cos\theta\langle\hat{\sigma}_{x}\rangle_{s}+\sin\theta\langle\hat{\sigma}_{z}\rangle_{s}$ and $\langle\hat{\sigma}_{z}\rangle=\cos\theta\langle\hat{\sigma}_{z}\rangle_{s}-\sin\theta\langle\hat{\sigma}_{x}\rangle_{s}$ of one parity-breaking solution. We set $r=10$ and $\theta=\frac{\pi}{7},\frac{2\pi}{7},\cdots,\frac{6\pi}{7}$. Other parameters are same as those in Fig. 2.
@@ III Exact numerical simulation
In the preceding section, we examined the metastability of the parity-breaking solutions using mean-field theory and cumulant expansion method. Although these methods provide evidence for the metastable nature of the SMP, they do not clarify its underlying physical origins. To address this limitation, we perform exact numerical simulations of the master equation shown in Eq. (2) for finite $\Omega/\omega_{0}$ and then extrapolate the results to $\Omega/\omega_{0}\rightarrow\infty$ through finite-size scaling analysis.
@@ III.1 Steady-state distribution
To further confirm the matestability of the parity-breaking solution in the SMP phase, we resort to exact numerical simulation of the system for finite $\Omega/\omega_{0}=1200$. By introducing a truncation of the Fock space of the bosonic mode, we numerically solve the master equation shown in Eq. (2). The steady state is obtained by solving the equation
@@
where $\mathcal{L}$ is defined as the Liouvillian of the system. The state of the bosonic mode is then obtained by a partial trace of the spin degrees of freedom, denoted as $\rho_{B}={\rm{Tr}}_{S}[\rho_{s}]$. In Fig. 4, we present the Husimi $Q$-representation of $\rho_{B}$ at different phases. The $Q$-representation is defined as
@@
where $|\alpha\rangle$ is the coherent state of the bosonic mode. It provides a quasi-probability distribution of the bosonic mode in phase space, offering a smooth and non-negative visualization of a quantum state.
In the NP, the $Q$-representation exhibits a single peak centered at the origin, as shown in Fig. 4(a), corresponding to the parity-preserving solution predicted by mean-field theory. The broadening of the peak is attributed to quantum fluctuations. As the coupling strength $\lambda$ increases beyond the critical point, the $Q$-representation develops two symmetric peaks away from the origin, as shown in Fig. 4(b), corresponding to the pair of parity-breaking solutions predicted by mean-field theory. The appearance of the paired peaks might suggest that the parity-breaking solutions are stable. However, a closer inspection of the $Q$-representation by rescaling the plot range reveals two distinct channels connecting the parity-breaking solutions and the saddle solution with finite, non-vanishing probability, as shown in Fig. 4(c). Specifically, a weak spiral channel provides a path that can drive the system from the parity-breaking solutions to the saddle point, whereas the sigmoid channel provides a path to settle the system into one of the parity-breaking solutions. The presence of these channels confirms that the parity-breaking solutions are metastable, consistent with the behavior revealed by the evolution of the symmetry-breaking solution under strong perturbations, as shown in Fig. 3(b). Consequently, steady state $\rho_{s}$ in the SMP retains the parity, causing $\langle\hat{x}\rangle_{s}=\langle\hat{p}\rangle_{s}=0$ and large quantum fluctuations. It is these enhanced quantum fluctuations that lead to the divergence of higher-order cumulants.
Figure 4: Husimi $Q$-representation of the bosonic mode in the NP (a) and SMP (b). Panel (c) shows a magnified view of (b), where the low-intensity regions are emphasized by rescaling the color map. Here $x\equiv\sqrt{2}\rm{Re}[\alpha]$ and $p\equiv\sqrt{2}\rm{Im}[\alpha]$. Parameters are same as those in Fig. 2.
@@ III.2 Principal component analysis
In the above analysis, we assume that there is a one-to-one correspondence between the steady-state distribution and the dynamics of the parity-breaking solutions. This correspondence arises from the distinctive nature of the steady state in an open quantum system, which is governed by a dynamical equilibrium among multiple contributing components. This stands in contrasts to the stationary state of a closed system, which is uniquely determined by the system’s Hamiltonian or free energy and typically consists of a single component. When the steady state contains only one single component, it coincides with the stationary state. However, when two or more components coexist, the steady state reflects the dynamical balance between them and thus faithfully reflects the evolution of these components.
@@
where $|\psi_{i}\rangle$ denotes the $i$-th eigenstate with probability $p_{i}$. Numerical results show that, in the SMP, two states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ dominant with equal probability, corresponding to the parity-breaking solutions, while other states only have quite small probabilities. A further calculation of the $Q$-representation of the bosonic mode reveals that each eigenstate exhibits certain features of the steady-state distribution shown in Fig. 4(c). See Appendix B for details. These features can be further classified into different groups according to their similarity. Such a grouping of the eigenstates yields a principal-component analysis of the steady state
@@
where $\rho^{c}_{l}$ is defined as the principal component $l$ and $g_{l}$ denotes the set of eigenstates in this group. Using the standard community detection algorithm developed for graph theory, we find four principal components in $\rho_{s}$. They are a pair of parity-breaking components, as shown in Fig. 5(a) and (b), one spiral component, and one sigmoid component, as shown in Fig. 5(c) and (d). Interpreting the parity-breaking components as metastable states and the spiral and sigmoid components as transient states, the principal component analysis establishes a correspondence between the steady-state distribution and the dynamics of the metastable states shown in Sect. II.3.
Figure 5: Principal components of the steady state in the SMP. Panels (a) and (b) reveal two parity-breaking components, while (c) and (d) show a spiral and a sigmoid components, respectively. Parameters are the same as those in Fig. 2.
@@
Here $0\leq F(P_{\psi_{i}},P_{\psi_{j}})\leq 1$ with $F=1$ when $P_{\psi_{i}}=P_{\psi_{j}}$ and $F=0$ when $P_{\psi_{i}}$ and $P_{\psi_{j}}$ do no overlap. We construct a weighted similarity graph of eigenstates based on their similarity, where each eigenstate is represented as a node, and two nodes are connected by an edge weighted by their similarity whenever it exceeds a chosen threshold [Newman2018]. Applying a standard community detection algorithm to this network with an edge threshold of $0.3$ yields the four principal components shown in Fig. 5.
The principal component analysis not only reveals the different components in the steady state but also provides an estimation of the relative lifetime of these components. Owing to one-to-one correspondence between the steady-state distribution and the dynamic balance among different components, the relative probabilities directly reflect their relative lifetimes, i.e.,
@@
Numerical results show that the probability of the parity-breaking component is significantly larger than those of the spiral and sigmoid components. Accordingly, the parity-breaking component has a longer lifetime and can thus be regarded as metastable states, whereas the spiral and sigmoid components are transient. A detail analysis of the spin probability distribution further reveals that the spin in the spiral component is dominated by the $|+\rangle_{z}$ state, whereas in the sigmoid component it is dominated by the $|-\rangle_{z}$ state (see Fig. 6(a) for details). Hence, the transition from the spiral component to the sigmoid component is associated with a spin jump, which yields the lifetime $1/2\gamma$ of each spiral channel. Further taking into account the two spiral channels in the spiral component, the lifetime of the parity-breaking state can then be estimated as
@@ III.3 Quantum trajectory simulation
Figure 6: Quantum trajectory of the parity-breaking state $|\psi_{1}\rangle$ obtained via the quantum jump method. (a) Time evolution of $\langle\hat{\sigma}_{z}(t)\rangle$, $\langle\hat{x}(t)\rangle$, and the fluctuation $\Delta x(t)=\sqrt{\langle\hat{x}^{2}(t)\rangle-\langle\hat{x}(t)\rangle^{2}}$. Dashed lines denotes the time at which the state jumps to the saddle point. Insets display the $Q$-representation of the bosonic mode before and after the saddle point at $\omega_{0}t=150.0$ and $\omega_{0}t=400.0$, respectively. (b) Evolution path of the $\{x(t)$, p(t)p(t)} (thick lines) reconstructed by tracking the positions of the peaks in the $Q$-representation of the trajectory state. Parameters are the same as those in Fig. 2.
To gain deeper insight into the correspondence between dynamics of metastable states and the steady-state distribution, we investigate the system dynamics from an initial parity-breaking state using the quantum jump method [Dalibard1992, Dum1992, Plenio1998, Link2019]. In this method, the system state $|\psi(t)\rangle$ evolves unitarily under the Hamiltonian $\hat{H}$, but is interrupted by stochastic quantum jumps induced by the bosonic mode relaxation $\mathcal{D}_{\hat{a}}$ or the spin relaxation $\mathcal{D}_{\hat{\sigma}_{-}}$. The state $|\psi(t)\rangle$ after a small time step $\delta t$ is updated according to
@@
Taking a representative trajectory as an example, we illustrate the evolution of the parity-breaking state under quantum jumps in the SMP. Starting from a symmetry-breaking state $|\psi_{1}\rangle$ defined in Eq. (14), the quantum trajectory is shown in Fig. 6 by taking $\langle\hat{\sigma}_{z}(t)\rangle$, $\langle\hat{x}(t)\rangle$, and $\Delta x(t)\equiv\sqrt{\langle\hat{x}^{2}(t)\rangle-\langle\hat{x}(t)\rangle^{2}}$ as figures of merit. We find that, in the absence of spin jumps (i.e., sharp spikes in $\langle\hat{\sigma}_{z}(t)\rangle$), the parity-breaking state remains stable. Each spin jump introduces a strong perturbation to the system, manifesting as a sharp spike in $\langle\hat{\sigma}_{z}(t)\rangle$. However, in most cases, these perturbations cannot drive the system away from the parity-breaking state, as evidenced by the unaffected evolution of $\langle\hat{x}(t)\rangle$, indicating the robustness of the parity-breaking state. Nevertheless, a small fraction of spin jumps can perturb the system to a transient state with $\langle\hat{\sigma}_{z}(t)\rangle\approx 1$ and an oscillatory decay of $\langle\hat{x}(t)\rangle$, thereby explicitly demonstrating the metastability of the parity-breaking state. In this case, the system evolves along one of the spiral channel depicted in Fig. 5(c). A subsequent spin jump (indicated by the dashed line) brings the system to the saddle point characterized by $\langle\hat{\sigma}_{z}\rangle\approx-1$ and $\langle\hat{x}(t)\rangle\approx 0$. Since the saddle point is unstable, small quantum fluctuations would further cause the system to rapidly relax back to one parity-breaking solution along the sigmoid-shaped channel depicted in Fig. 5(d). Consequently, the trajectory simulation provides direct evidence of the metastable nature of the parity-breaking state.
@@
Notably, the relaxation from the saddle point back to the parity-breaking state is not directly visible in $\langle\hat{x}(t)\rangle$. This is because the parity symmetry is restored at the saddle point, and the subsequent dynamics governed by the quantum jump equation Eq. (20) cannot break this symmetry. Consequently, $\langle\hat{\sigma}_{x}(t)\rangle=0$ holds throughout the ensuing dynamics. In this regime, the trajectory state $|\psi(t)\rangle$ forms a cat state–a superposition of two distinct parity-breaking states–manifested by large fluctuations in $\hat{x}(t)$. This behavior is evident in the evolution of $\Delta x(t)$, which shows a sharp increase after the jump to the saddle point (denoted by the dashed line). The $Q$-representation of the bosonic mode before and after the saddle point, shown in the insets, further illustrates the emergence of the cat state. To obtain the evolution path of the system, we identify the peak in the $Q$-representation and use its position $\{x(t),p(t)\}$ as the representative coordinate. The reconstructed evolution path $\{x(t),p(t)\}$ is shown as thick lines in Fig. 6(b). A bifurcation of the evolution path (highlighted by blue lines) emerges due to the cat state nature of the $|\psi(t)\rangle$. The evolution path establishes a one-to-one correspondence between the dynamical evolution of the parity-breaking state and the steady-state distribution, thereby confirming the conclusion in Sect. III.2 and the metastability of the parity-breaking state.

### v2

@@ II.3 Stability under strong perturbations
Figure 3: Evolution of the steady-state solution in the (a) NP and (b) SMP under strong perturbations within the mean-field framework. Phenomenological interpretations are shown in the top panels. For the NP, we consider a quench of the parity-preserving solution with $\langle\hat{x}\rangle=r\cos 2\theta$ and $\langle\hat{p}\rangle=r\sin 2\theta$. For the SMP, we consider a quench of one parity-breaking solution with $\langle\hat{\sigma}_{x}\rangle=\cos\theta\langle\hat{\sigma}_{x}\rangle_{s}+\sin\theta\langle\hat{\sigma}_{z}\rangle_{s}$ and $\langle\hat{\sigma}_{z}\rangle=\cos\theta\langle\hat{\sigma}_{z}\rangle_{s}-\sin\theta\langle\hat{\sigma}_{x}\rangle_{s}$. We set $r=10$ and $\theta=\frac{\pi}{7},\frac{2\pi}{7},\cdots,\frac{6\pi}{7}$. The other parameters are the same as those in Fig. 2.
@@
From the cumulant analysis, we have shown that strong quantum fluctuations exist in the SMP. This indicates that the parity-breaking solutions in the SMP can be stable only for a finite time before a large perturbation drives the system away. We thus refer to these parity-breaking solutions as metastable solutions and label the phase as a superradiant metastable phase.
@@ III Exact numerical simulation
In the preceding section we examined the metastability of the parity-breaking solutions using mean-field theory and the cumulant expansion method. Although these methods provide evidence for the metastable nature of the SMP, they do not clarify its underlying physical origins. To address this limitation, we perform exact numerical simulations of the master equation (2) for finite $\Omega/\omega_{0}$ and then extrapolate the results to $\Omega/\omega_{0}\rightarrow\infty$ through finite-size scaling analysis.
@@ III.1 Steady-state distribution
To further confirm the metastability of the parity-breaking solution in the SMP phase, we resort to exact numerical simulation of the system for finite $\Omega/\omega_{0}=1200$. By introducing a truncation of the Fock space of the bosonic mode, we numerically solve the master equation (2). The steady state is obtained by solving the equation
@@
where $\mathcal{L}$ is defined as the Liouvillian of the system. The state of the bosonic mode is then obtained by a partial trace of the spin degrees of freedom, defined as $\rho_{B}={\rm{Tr}}_{S}[\rho_{s}]$. In Fig. 4 we present the Husimi $Q$ representation of $\rho_{B}$ at different phases. The $Q$ representation is defined as
@@
where $|\alpha\rangle$ is the coherent state of the bosonic mode. It provides a quasiprobability distribution of the bosonic mode in phase space, offering a smooth and non-negative visualization of a quantum state.
In the NP, the $Q$ representation exhibits a single peak centered at the origin, as shown in Fig. 4(a), corresponding to the parity-preserving solution predicted by mean-field theory. The broadening of the peak is attributed to quantum fluctuations. As the coupling strength $\lambda$ increases beyond the critical point, the $Q$ representation develops two symmetric peaks away from the origin, as shown in Fig. 4(b), corresponding to the pair of parity-breaking solutions predicted by mean-field theory. The appearance of the paired peaks might suggest that the parity-breaking solutions are stable. However, a closer inspection of the $Q$ representation by rescaling the plot range reveals two distinct channels connecting the parity-breaking solutions and the saddle solution with finite nonvanishing probability, as shown in Fig. 4(c). Specifically, a weak spiral channel provides a path that can drive the system from the parity-breaking solutions to the saddle point, whereas the sigmoid channel provides a path to settle the system into one of the parity-breaking solutions. The presence of these channels confirms that the parity-breaking solutions are metastable, consistent with the behavior revealed by the evolution of the symmetry-breaking solution under strong perturbations, as shown in Fig. 3(b). Consequently, steady state $\rho_{s}$ in the SMP retains the parity, causing $\langle\hat{x}\rangle_{s}=\langle\hat{p}\rangle_{s}=0$ and large quantum fluctuations. It is these enhanced quantum fluctuations that lead to the divergence of higher-order cumulants.
@@ III.2 Principal component analysis
In the above analysis we assumed that there is a one-to-one correspondence between the steady-state distribution and the dynamics of the parity-breaking solutions. This correspondence arises from the distinctive nature of the steady state in an open quantum system, which is governed by a dynamical equilibrium among multiple contributing components. This stands in contrast to the stationary state of a closed system, which is uniquely determined by the system’s Hamiltonian or free energy and typically consists of a single component. When the steady state contains only one single component, it coincides with the stationary state. However, when two or more components coexist, the steady state reflects the dynamical balance between them and thus faithfully reflects the evolution of these components.
@@
where $|\psi_{i}\rangle$ denotes the $i$th eigenstate with probability $p_{i}$. Numerical results show that, in the SMP, two states $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ dominant with equal probability, corresponding to the parity-breaking solutions, while other states only have quite small probabilities. A further calculation of the $Q$ representation of the bosonic mode reveals that each eigenstate exhibits certain features of the steady-state distribution shown in Fig. 4(c) (see Appendix B for details). These features can be further classified into different groups according to their similarity. Such a grouping of the eigenstates yields a principal-component analysis of the steady state
@@
where $\rho^{c}_{l}$ is defined as the principal component $l$ and $g_{l}$ denotes the set of eigenstates in this group. Using the standard community detection algorithm developed for graph theory, we find four principal components in $\rho_{s}$. They are a pair of parity-breaking components, as shown in Figs. 5(a) and (b), one spiral component [Fig. 5(c)], and one sigmoid component [Fig. 5(d)]. Interpreting the parity-breaking components as metastable states and the spiral and sigmoid components as transient states, the principal component analysis establishes a correspondence between the steady-state distribution and the dynamics of the metastable states shown in Sec. II.3.
@@
Here $0\leq F(P_{\psi_{i}},P_{\psi_{j}})\leq 1$ with $F=1$ when $P_{\psi_{i}}=P_{\psi_{j}}$ and $F=0$ when $P_{\psi_{i}}$ and $P_{\psi_{j}}$ do no overlap. We construct a weighted similarity graph of eigenstates based on their similarity, where each eigenstate is represented as a node, and two nodes are connected by an edge weighted by their similarity whenever it exceeds a chosen threshold [52]. Applying a standard community detection algorithm to this network with an edge threshold of $0.3$ yields the four principal components shown in Fig. 5.
Figure 4: Husimi $Q$ representation of the bosonic mode in the (a) NP and (b) SMP. (c) Magnified view of (b), where the low-intensity regions are emphasized by rescaling the colormap. Here $x\equiv\sqrt{2}\rm{Re}[\alpha]$ and $p\equiv\sqrt{2}\rm{Im}[\alpha]$. The parameters are the same as those in Fig. 2.
The principal component analysis not only reveals the different components in the steady state but also provides an estimation of the relative lifetime of these components. Owing to the one-to-one correspondence between the steady-state distribution and the dynamic balance between different components, the relative probabilities directly reflect their relative lifetimes, i.e.,
@@
Numerical results show that the probability of the parity-breaking component is significantly larger than those of the spiral and sigmoid components. Accordingly, the parity-breaking component has a longer lifetime and can thus be regarded as a metastable state, whereas the spiral and sigmoid components are transient. A detailed analysis of the spin probability distribution further reveals that the spin in the spiral component is dominated by the $|+\rangle_{z}$ state, whereas in the sigmoid component it is dominated by the $|-\rangle_{z}$ state [see Fig. 6(a) for details]. Hence, the transition from the spiral component to the sigmoid component is associated with a spin jump, which yields the lifetime $1/2\gamma$ of each spiral channel. Further taking into account the two spiral channels in the spiral component, the lifetime of the parity-breaking state can then be estimated as
@@
Figure 5: Principal components of the steady state in the SMP. (a) and (b) two parity-breaking components, (c) a spiral and a sigmoid component, and (d) show a sigmoid component. The parameters are the same as those in Fig. 2.
@@ III.3 Quantum trajectory simulation
Figure 6: Quantum trajectory of the parity-breaking state $|\psi_{1}\rangle$ obtained via the quantum jump method. (a) Time evolution of $\langle\hat{\sigma}_{z}(t)\rangle$, $\langle\hat{x}(t)\rangle$, and the fluctuation $\Delta x(t)=\sqrt{\langle\hat{x}^{2}(t)\rangle-\langle\hat{x}(t)\rangle^{2}}$. The vertical dotted line denotes the time at which the state jumps to the saddle point. Insets display the $Q$ representation of the bosonic mode before and after the saddle point at $\omega_{0}t=150.0$ and $\omega_{0}t=400.0$, respectively. (b) Evolution path of $\{x(t)$, p(t)p(t)} (thick lines) reconstructed by tracking the positions of the peaks in the $Q$ representation of the trajectory state. The parameters are the same as those in Fig. 2.
To gain deeper insight into the correspondence between the dynamics of metastable states and the steady-state distribution, we investigate the system dynamics from an initial parity-breaking state using the quantum jump method [14, 16, 55, 40]. In this method, the system state $|\psi(t)\rangle$ evolves unitarily under the Hamiltonian $\hat{H}$, but is interrupted by stochastic quantum jumps induced by the bosonic mode relaxation $\mathcal{D}_{\hat{a}}$ or the spin relaxation $\mathcal{D}_{\hat{\sigma}_{-}}$. The state $|\psi(t)\rangle$ after a small time step $\delta t$ is updated according to
@@
Taking a representative trajectory as an example, we illustrate the evolution of the parity-breaking state under quantum jumps in the SMP. Starting from a symmetry-breaking state $|\psi_{1}\rangle$ defined in Eq. (14), the quantum trajectory is shown in Fig. 6 by taking $\langle\hat{\sigma}_{z}(t)\rangle$, $\langle\hat{x}(t)\rangle$, and $\Delta x(t)\equiv\sqrt{\langle\hat{x}^{2}(t)\rangle-\langle\hat{x}(t)\rangle^{2}}$ as figures of merit. We find that, in the absence of spin jumps, i.e., sharp spikes in $\langle\hat{\sigma}_{z}(t)\rangle$, the parity-breaking state remains stable. Each spin jump introduces a strong perturbation to the system, manifesting as a sharp spike in $\langle\hat{\sigma}_{z}(t)\rangle$. However, in most cases, these perturbations cannot drive the system away from the parity-breaking state, as evidenced by the unaffected evolution of $\langle\hat{x}(t)\rangle$, indicating the robustness of the parity-breaking state. Nevertheless, a small fraction of spin jumps can perturb the system to a transient state with $\langle\hat{\sigma}_{z}(t)\rangle\approx 1$ and an oscillatory decay of $\langle\hat{x}(t)\rangle$, thereby explicitly demonstrating the metastability of the parity-breaking state. In this case, the system evolves along one of the spiral channel depicted in Fig. 5(c). A subsequent spin jump (indicated by the vertical dotted line) brings the system to the saddle point characterized by $\langle\hat{\sigma}_{z}\rangle\approx-1$ and $\langle\hat{x}(t)\rangle\approx 0$. Since the saddle point is unstable, small quantum fluctuations would further cause the system to rapidly relax back to one parity-breaking solution along the sigmoid-shaped channel depicted in Fig. 5(d). Consequently, the trajectory simulation provides direct evidence of the metastable nature of the parity-breaking state.
@@
Notably, the relaxation from the saddle point back to the parity-breaking state is not directly visible in $\langle\hat{x}(t)\rangle$. This is because the parity symmetry is restored at the saddle point, and the subsequent dynamics governed by the quantum jump equation (20) cannot break this symmetry. Consequently, $\langle\hat{\sigma}_{x}(t)\rangle=0$ holds throughout the ensuing dynamics. In this regime, the trajectory state $|\psi(t)\rangle$ forms a cat state, i.e., a superposition of two distinct parity-breaking states, manifested by large fluctuations in $\hat{x}(t)$. This behavior is evident in the evolution of $\Delta x(t)$, which shows a sharp increase after the jump to the saddle point (denoted by the dotted line). The $Q$ representation of the bosonic mode before and after the saddle point, shown in the insets, further illustrates the emergence of the cat state. To obtain the evolution path of the system, we identify the peak in the $Q$ representation and use its position $\{x(t),p(t)\}$ as the representative coordinate. The reconstructed evolution path $\{x(t),p(t)\}$ is shown as thick lines in Fig. 6(b). A bifurcation of the evolution path (highlighted by blue lines) emerges due to the cat state nature of the $|\psi(t)\rangle$. The evolution path establishes a one-to-one correspondence between the dynamical evolution of the parity-breaking state and the steady-state distribution, thereby confirming the conclusion in Sec. III.2 and the metastability of the parity-breaking state.